A certain right triangle has its area numerically equal to its perimeter. The length of each side is an even integer. What is the perimeter?
(1) 26 (2) 28 (3) 30 (4) 32 (5) None of the above

Found 3 solutions by palanisamy, richwmiller, Alan3354:
Answer by palanisamy(496)

The sides of a right triangle are in the ratio 3:4:5
Since the sides are even numbers we can take the sides to be 6,8,10
Perimeter = 6+8+10 = 24
Area = (1/2)*6*8 = 24
Therefore the required perimeter is 24

Answer by richwmiller(17219)

Area=1/2 bh
p=b+h+c
p=6+8+10=24
area=1/2*6*8=24
none of the above

Answer by Alan3354(69443)

The sides are 5, 12 and 13.
Area = 5*12/2 = 30
P = 5+12+13 = 30